# The Fellowship of the LLMs: Multi-Agent Workflows for Synthetic Preference Optimization Dataset Generation

### v4

@@ 4.2 LLM-as-Generator
We compare the performance of Multi-Agent Feedback Loop with the baseline Single-Agents (GPT-4o, Llama-3.1-8b, Gemma-2-9b) using win rate as shown in Table 6.
@@
Table 6: Win Rate of Multi-Agent and Single-Agent against GPT-4o, Llama-3.1-8b and Gemma-2-9b
@@
In the Multi-Agent setting, all variations outperform the single-agents against GPT-4o, with the highest win rate of 49.0% for Llama as a generator and Gemma as a reviewer. This configuration performs the best against Llama and Gemma too, with 71.8% and 73.8% win rate respectively. We observe that using Llama as the generator improves the performance as compared to using Gemma as the generator because this configuration leads to a better win rate against all three baselines.
Llama’s strengths in generating responses may be enhanced by Gemma’s ability to fine-tune and correct the errors, leading to more polished outputs. The results underscore the importance of assigning appropriate roles based on the specific strengths of each model. Llama, when set as the generator, appears to leverage its capabilities more effectively than Gemma in this role. The use of diverse models in the feedback loop likely helps mitigate biases that any single model might introduce. This diversity ensures a broader range of perspectives while answer a question. In conclusion, the demonstrated efficacy of the Multi-Agent Feedback Loop, especially with Llama as the generator and Gemma as the reviewer, validates the concept of collaborative AI systems.
@@ 4.3 Preference Optimization Dataset
Figure 5: Multi-agent framework for PO dataset generation.
@@ 5 Conclusion
This paper presents PO datasets generated using multi-agent frameworks, and evaluates these frameworks by highlighting the advantages, drawbacks, and challenges of each approach. In the response evaluation module, our comparative analysis of LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate shows the suitability of each setup depending on the context of use. For the response generation module, we evaluate the LLM Feedback Loop using Llama-3.1-8b and Gemma-2-9b in various configurations. LLM-as-a-Judge proved to be highly effective when candidate responses don’t have a response from the Judge LLM. Whereas LLMs-as-a-Jury and LLM Debate demonstrated robustness, particularly useful in reducing evaluator bias. However, Cohen’s Kappa for both of these approaches has a high variance making them less suitable for novel applications.
Our experiments with LLM Feedback Loop using Llama-3.1-8b and Gemma-2-9b configurations show the potential of multi-agent frameworks in refined content generation. Configurations where Llama-3.1-8b served as the generator and Gemma-2-9b as the reviewer consistently delivered better results, demonstrating the benefits of leveraging complementary strengths of different models to refine output quality. These findings indicate the effectiveness of multi-agent frameworks for varied AI applications, showing promise for moving towards systems requiring minimal human intervention - however, this method is computationally expensive in comparison.
We also generate multiple DPO and KPO datasets using LLM Feedback Loop with Llama-3.1-8b as the generator and Gemma-2-9b as the evaluator and GPT-4o-as-a-Judge. The aim of these datasets is to improve single-agent capabilities for better response generation and multi-agent capabilities including better communication and improved feedback.
In order to facilitate further research and ensure transparency, all code, LLM responses, and generated datasets have been made public.
@@ 6 Future Work
7 Limitations
While our study demonstrates the potential of multi-agent workflows in automating the generation of PO datasets, several limitations should be acknowledged. Firstly, the use of multi-agent frameworks significantly increases computational complexity and resource consumption compared to single-agent models. The iterative processes in both the response generation and evaluation modules require more computational power and time, which may not be feasible for practitioners with limited resources. Additionally, GPT-4o is a proprietary model, which may not be accessible to all researchers, potentially hindering reproducibility and wider adoption of our methods.
8 Ethical Considerations
@@
Acknowledgments
We are grateful to OpenAI for supporting our work through their Research Access Program.
@@ Appendix B Code and Datasets
The evaluation code, all the evaluation outputs and the generated datasets are publicly available on GitHub [^4] https://github.com/ulrs0/MA-PO. For evaluation of LLMs-as-Evaluators we used Alpaca-Eval [^5] https://huggingface.co/datasets/tatsu-lab/alpaca_eval, Fair-Eval [^6] https://github.com/i-Eval/FairEval, PandaLM-Eval [^7] https://github.com/WeOpenML/PandaLM and MT-Bench [^8] https://huggingface.co/datasets/lmsys/mt_bench_human_judgments. For evaluation of LLMs-as-Generators and single-agent improvement dataset generation we use the prompts from Argilla Capybara DPO dataset [^9] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized. For multi-agent improvement dataset generation we use prompts from No-Robots [^10] https://huggingface.co/datasets/HuggingFaceH4/no_robots dataset. Alpaca-Eval and PandaLM-Eval are under Apache 2.0 license, Fair-Eval dataset is under CC BY 4.0 license, Argilla Capybara DPO is also under Apache 2.0 license. All datasets used in this paper comply with their respective license.

### v5

@@ 4.2 LLM-as-Generator
We compare the performance of multi-model Feedback Loop with the baseline single-models (GPT-4o, Llama-3.1-8b, Gemma-2-9b) using win rate as shown in Table 7.
@@
Table 7: Win Rate of multi-model and single-model against GPT-4o, Llama-3.1-8b and Gemma-2-9b
@@
In the multi-model setting, all variations outperform the single-models against GPT-4o, with the highest win rate of 49.0% for Llama as a generator and Gemma as a reviewer. This configuration performs the best against Llama and Gemma too, with 71.8% and 73.8% win rate respectively. We observe that using Llama as the generator improves the performance as compared to using Gemma as the generator because this configuration leads to a better win rate against all three baselines.
Llama’s strengths in generating responses may be enhanced by Gemma’s ability to fine-tune and correct the errors, leading to more polished outputs. The results underscore the importance of assigning appropriate roles based on the specific strengths of each model. Llama, when set as the generator, appears to leverage its capabilities more effectively than Gemma in this role. The use of diverse models in the feedback loop likely helps mitigate biases that any single model might introduce. This diversity ensures a broader range of perspectives while answer a question. In conclusion, the demonstrated efficacy of the multi-model Feedback Loop, especially with Llama as the generator and Gemma as the reviewer, validates the concept of collaborative AI systems.
@@ 4.3 Preference Optimization Dataset
Figure 4: Multi-model framework for PO dataset generation.
@@ 5 Conclusion
This paper presents PO datasets generated using multi-model frameworks, and evaluates these frameworks by highlighting the advantages, drawbacks, and challenges of each approach. In the response evaluation module, our comparative analysis of LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate shows the suitability of each setup depending on the context of use. For the response generation module, we evaluate the LLM Feedback Loop using Llama-3.1-8b and Gemma-2-9b in various configurations. LLM-as-a-Judge proved to be highly effective when candidate responses don’t have a response from the Judge LLM. Whereas LLMs-as-a-Jury and LLM Debate demonstrated robustness, particularly useful in reducing evaluator bias. However, Cohen’s Kappa for both of these approaches has a high variance making them less suitable for novel applications.
Our experiments with LLM Feedback Loop using Llama-3.1-8b and Gemma-2-9b configurations show the potential of multi-model frameworks in refined content generation. Configurations where Llama-3.1-8b served as the generator and Gemma-2-9b as the reviewer consistently delivered better results, demonstrating the benefits of leveraging complementary strengths of different models to refine output quality. These findings indicate the effectiveness of multi-model frameworks for varied AI applications, showing promise for moving towards systems requiring minimal human intervention - however, this method is computationally expensive in comparison.
We also generate multiple DPO and KPO datasets using LLM Feedback Loop with Llama-3.1-8b as the generator and Gemma-2-9b as the evaluator and GPT-4o-as-a-Judge. The aim of these datasets is to improve single-model capabilities for better response generation and multi-model capabilities including better communication and improved feedback.
@@ 6 Future Work
Limitations
While our study demonstrates the potential of multi-model workflows in automating the generation of PO datasets, several limitations should be acknowledged. Firstly, the use of multi-model frameworks significantly increases computational complexity and resource consumption compared to single-model models. The iterative processes in both the response generation and evaluation modules require more computational power and time, which may not be feasible for practitioners with limited resources. Additionally, GPT-4o is a proprietary model, which may not be accessible to all researchers, potentially hindering reproducibility and wider adoption of our methods.
Ethical Considerations
@@ Appendix B Code and Datasets
The evaluation code, all the evaluation outputs and the generated datasets are publicly available on GitHub [^4] https://github.com/sameearif/Fellowship-of-The-LLMs. For evaluation of LLMs-as-Evaluators we used Alpaca-Eval [^5] https://huggingface.co/datasets/tatsu-lab/alpaca_eval, Fair-Eval [^6] https://github.com/i-Eval/FairEval, PandaLM-Eval [^7] https://github.com/WeOpenML/PandaLM and MT-Bench [^8] https://huggingface.co/datasets/lmsys/mt_bench_human_judgments. For evaluation of LLMs-as-Generators and single-model improvement dataset generation we use the prompts from Argilla Capybara DPO dataset [^9] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized. For multi-model improvement dataset generation we use prompts from No-Robots [^10] https://huggingface.co/datasets/HuggingFaceH4/no_robots dataset. Alpaca-Eval and PandaLM-Eval are under Apache 2.0 license, Fair-Eval dataset is under CC BY 4.0 license, Argilla Capybara DPO is also under Apache 2.0 license. All datasets used in this paper comply with their respective license.